# The Psychology of Learning from Machines: Anthropomorphic AI and the Paradox of Automation in Education

### v1

@@ I Introduction
Recent surveys indicate rapid increases in instructor adoption of AI tutoring systems, with the majority of students now regularly using AI tools for learning support [genai_adoption_stats]. This widespread deployment raises a fundamental question about the nature of these systems and our interactions with them. We must distinguish between two distinct modes of AI-mediated learning: conversing with conversational AI agents as virtual partners versus learning from artifacts created by generative AI (GenAI). The former involves direct dialogue with systems that simulate human-like social presence, while the latter treats AI as an intelligent tool whose outputs we consume and evaluate. This distinction matters because each mode evokes different psychological responses, ethical considerations, and pedagogical implications.
@@
This difference connects to broader philosophical debates about technology’s role in human activity. Are AI tutors instruments—tools that help us reach specific educational ends—or are they becoming agents with whom we interact socially and to whom we may inappropriately attribute understanding, care, or authority? The philosophy of technology offers multiple lenses: technological determinism suggests technology shapes society independent of human will; technological instrumentalism views technology as neutral means to human-chosen ends; technological co-evolution recognizes mutual shaping between humans and artifacts [winograd1987understanding]. Modern AI tutors occupy an ambiguous position across this spectrum—their conversational fluency and apparent understanding create what Sedlakova and Trachsel call a “hybrid nature,” evoking familiar categories of human interaction while not truly corresponding to them [sedlakova2023conversational].
@@
This ambiguity creates what we term the personhood conferral problem: when AI agents imitate not just explanation but human presence, they can begin to assume roles traditionally reserved for human educators—mentor, evaluator, emotional supporter, intellectual authority [stein2024personhood]. The tendency to attribute human-like qualities to non-human entities is ancient; In the sixth century BC, Xenophanes critiqued humans for imagining gods in their own image—an early articulation of what we now call anthropomorphism—and centuries later Hume described a “universal tendency among mankind to conceive all beings like themselves.” Yet modern large language models (LLMs) amplify this tendency through unprecedented conversational fluency, producing contextually appropriate responses that create compelling illusions of genuine comprehension [reeves1996media, epley2007seeing, xiao2025humanizing].
@@
The generative capacity of anthropomorphic AI has both pedagogical promise and profound risks [qadir2023engineering, qadir2024educating]. AI offers personalized feedback, adaptive scaffolding, and simulated dialogues that support complex reasoning; yet these same affordances blur boundaries between authentic and synthetic learning, raising concerns about trust calibration, cognitive dependency, epistemic capture, and the preservation of distinctively human elements of education [holmes2025critical, tabassum2024critical]. The mutual relationship between anthropomorphism of technology and dehumanization of educational processes deserves careful examination—as we make machines more human-like, we risk making human processes more mechanical.
@@
This paper addresses a critical question: What can established research reveal about the psychology of learning from anthropomorphic AI systems in educational contexts? Educational institutions are integrating AI systems faster than research can evaluate their psychological and pedagogical implications, creating a temporal paradox where deployment outpaces understanding. Yet this challenge is not entirely unprecedented. We possess decades of relevant research from human factors engineering, automation psychology, and human–computer interaction (HCI) examining how humans relate to intelligent machines across high-stakes domains from aviation to industrial control [bainbridge1983ironies, skitka1999automation, carr2014glass].
@@
Building on four decades of automation research [bainbridge1983ironies, simkute2025genai], our study integrates insights from multiple traditions—human factors engineering, automation psychology, HCI, educational technology, and philosophical analyses of machine limitations [winograd1987understanding, polanyi1966tacit]. Through theoretical synthesis combined with exploratory analysis of 72,178 naturalistic learner comments on AI-generated educational content, we identify three persistent psychological challenges:
@@
  • Trust Calibration Paradox: Users simultaneously overtrust AI through automation bias [skitka1999automation] and undertrust it through algorithm aversion [dietvorst2015algorithm], struggling to achieve appropriate reliance.
@@
  • Anthropomorphism Paradox: Users attribute human-like qualities to conversational agents despite knowing they are non-sentient [reeves1996media, epley2007seeing]. Design features increasing engagement can paradoxically distract from learning [bettysbrain] while creating risks of emotional dependency and inappropriate authority attribution.
@@
  • Automation Ironies Persist: Systems designed to support cognitive work inadvertently introduce new errors and degrade the very skills they aim to develop—a pattern documented in aviation [bainbridge1983ironies] now manifesting in education with GenAI [simkute2025genai, wharton_shallow].
@@
  • Design Principles and Future Research Directions: We translate these insights into guidelines for responsible educational AI design—highlighting transparency, scaffolding, learner agency, and cognitive load management—while identifying the personhood conferral problem as a central risk. We outline critical gaps requiring further study, including longitudinal impacts, cross-cultural variation, and domain-specific norms.
@@ II-A The Tool-Agent Spectrum and Hybrid Nature of AI
Technologies can be understood as neutral instruments extending human capability, as autonomous forces shaping society, or through frameworks recognizing mutual shaping between humans and artifacts [winograd1987understanding]. Modern AI tutors occupy ambiguous positions across the tool-agent spectrum. They function instrumentally when delivering content or checking answers, yet their conversational interfaces and apparent understanding invite agent-like attributions of intention, comprehension, and care. Unlike traditional tools serving purely instrumental purposes or human teachers possessing actual understanding and agency, AI tutors exist in liminal space: anthropomorphic enough to trigger social responses yet.
Sedlakova and Trachsel characterize chatbots as possessing a “hybrid nature”—entities that evoke familiar categories of personhood without genuinely instantiating them [sedlakova2023conversational]. fundamentally computational. Our lack of evolutionary experience with such ontologically confounding entities makes clear-headed evaluation exceptionally difficult, creating what we term the category confusion problem—users constantly navigate between treating AI as tool, partner, or oracle, often shifting framings within single interactions.
This creates the personhood conferral problem: explicit or implicit attribution of person-like status to AI systems [stein2024personhood]. When users interact with AI tutors using social scripts appropriate for human teachers—expressing gratitude, apologizing for mistakes, describing relationships—they may begin treating these systems as legitimate authorities, confidants, or evaluators. Granting personhood status to AI, either explicitly in legal code or implicitly in common practice, creates profound risks. This conferral carries multiple concerns: cognitive atrophy in foundational skills learned only through struggle with human feedback; epistemic capture where students overestimate AI reliability and intent; and emotional dependency displacing the relational complexity essential for professional identity formation.
@@ II-B Automation Psychology: Trust Calibration Challenges
Automation psychology has long grappled with paradoxes of human–machine interaction. Bainbridge’s seminal “Ironies of Automation” (1983) [bainbridge1983ironies] revealed a central irony persisting today: as systems become more automated, the need for skilled human judgment does not diminish—it grows. Drawing on aviation studies, she showed automation designed to eliminate human error often introduced designer errors, weakened operators’ skills through disuse, and reduced humans to passive monitoring—a task we perform poorly.
Psychologists documented how humans calibrate trust in automated systems. Skitka and colleagues [skitka1999automation] demonstrated operators using “very reliable but not perfectly reliable” automated aids performed worse than those without automation, committing omission errors (missing events when automation failed) at 55% rates and commission errors (following incorrect automation) approaching 100%. Reviews found even 80-90% accurate AI systems introduce new errors through overreliance, with users 26% more likely to make incorrect decisions when following erroneous AI advice [automation_bias_review].
Paradoxically, the reverse also occurs. Dietvorst and colleagues identified “algorithm aversion”—the tendency to lose confidence in algorithms more quickly than humans after observing identical mistakes [dietvorst2015algorithm]. This creates persistent trust calibration challenges: users may overtrust through uncritical acceptance or undertrust by rejecting systems prematurely after minor errors. The “expertise paradox” complicates this further—novices lacking confidence exhibit automation bias, while experts confident in their judgment show algorithm aversion [expertise_paradox].
In educational contexts, these patterns manifest distinctly. Students accept incorrect AI code recommendations without evaluation [programming_automation_bias]; mathematics learners show AI dependency correlating with reduced problem-solving [math_dependency]. A Wharton study reveals particularly insidious failure: students using GenAI performed better while practicing but 17% worse on exams when AI support was removed, compared with peers who never used AI [wharton_shallow]—dependency masked until assessment reveals skill deficits.
@@ II-C Anthropomorphism: Ancient Roots, Modern Amplification
Anthropomorphism—the tendency to attribute human qualities to non-human entities—has deep historical roots. Xenophanes observed that people imagined gods in their own image, and Hume later described a “universal tendency” to project human traits onto the world. What was once a cognitive habit grounded in myth and metaphor is now powerfully amplified by modern AI systems.
@@
Foundational studies such as the “Computers Are Social Actors” (CASA) paradigm showed that even minimal cues could trigger social responses toward computers [reeves1996media]. Although recent replications suggest these effects weaken with familiar technologies [casa_replication], contemporary conversational AI reactivates and intensifies them. Human-like names, voices, turn-taking behaviors, and chat interfaces are not neutral features: they are deliberate design choices that invite users to perceive understanding, agency, and relationality where none exists [gabriel2024ethics]. Similar patterns have been observed with digital voice assistants, whose polished personas lead users to overestimate their capabilities.
@@
Psychological theories deepen this picture. Epley et al.’s three-factor theory identifies the conditions that make anthropomorphism likely—pre-existing agent knowledge, the human need for social connection, and the desire for predictability and control [epley2007seeing]. Modern AI systems activate all three factors simultaneously. Xiao et al.’s Humanizing Machines framework [xiao2025humanizing] further shows that anthropomorphism is co-produced: designers embed perceptual, linguistic, and behavioral cues, and users complete the process through interpretation. They argue anthropomorphism should be viewed as a tunable design parameter—capable of supporting user goals when calibrated carefully, but hazardous when left unchecked.
@@
Linguistic cues alone can be powerful. DeVrio et al. [DeVrio2025Taxonomy] document nineteen linguistic constructions—from mental-state verbs to purposive expressions—that reliably evoke human-like agency. Even subtle phrasing can prompt users to attribute intention or understanding, especially when they are motivated to find an intelligent partner.
@@
These tendencies are not merely theoretical. Safety evaluations of GPT-4o report users expressing emotional attachment (e.g., “This is our last day together”) and warn that deferential, always-available AI could reshape norms of social interaction and diminish human-to-human connection [openai_gpt4o]. Human-like voice and persistent memory magnify trust miscalibration and the risk of over-reliance.
@@
In educational settings, the stakes are especially high. Students benefit when tutors appear attuned to their thoughts and emotions [tom_education], yet excessive sociability in AI tutors can distract and impair performance [bettysbrain]. More seriously, anthropomorphic attachment can lead to emotional dependency: Public Citizen’s 2023 report documents suicides linked to chatbot relationships [publiccitizen2023]. As the Ethics of Advanced AI Assistants report emphasizes, anthropomorphic features designed to enhance engagement carry significant ethical responsibility [gabriel2024ethics].
@@ II-D Anthropomorphism in the GenAI Era: Amplified Risks
The Ethics of Advanced AI Assistants report (Gabriel et al. [gabriel2024ethics]) highlights that advanced AI systems may be endowed with characteristics observed in social robots and digital assistants: embodiment, natural language interfaces, voice activation with realistic generation, and assertions of having identities, personalities, and internal states. These capabilities create new challenges for educational deployment requiring careful ethical consideration. Most critically, educational institutions deploy these systems faster than research can evaluate effects. This creates urgent need for both theoretical synthesis of existing knowledge and empirical analysis in real-world educational contexts. As Stein argues, “the risks from deeply anthropomorphic AI will begin to unfold in ways that are not obvious” [stein2024personhood]—how would we know when outcomes for youth inundated by technology reach critical failure such that reproduction of social roles needed for collective intelligence necessary for civilization does not occur? This makes it essential to leverage decades of relevant research while conducting contemporary analysis. Our work addresses this gap through systematic cross-disciplinary synthesis combined with large-scale exploratory analysis of naturalistic learner discourse.
@@ III-B YouTube Discourse Analysis
Analytical Procedures. We employed VADER (Valence Aware Dictionary and sEntiment Reasoner) [hutto2014vader] for polarity classification, validated for social-media text. Compound scores ranged from 1-1 (negative) to +1+1 (positive), categorized as positive ($\geq 0.05$), neutral (0.05<x<0.05-0.05<x<0.05), and negative ($\leq-0.05$). Temporal engagement was modeled using five logarithmically scaled bins: 0-1h (immediate), 1-6h (early), 6-24h (first day), 1-7d (first week), 7+d (long-tail). These bins balance early responsiveness with algorithmic discovery phases. Each YouTube comment was assigned to one of six engagement profiles using linguistic and structural features: (1) Analytical Inquirer (question + $\geq$2 analytical markers), (2) Critical Analyst ($\geq$3 analytical markers + $\geq$30 words), (3) Engaged Dialoguer ($\geq$50 words + analytical/emotional markers), (4) Emotional Broadcaster ($\geq$3 emotional markers or $\geq$100 words), (5) Casual Observer ($\leq$20 words), and (6) General Responder (default). Validation through cognitive-affective language ratio analysis confirmed a ninefold difference in reasoning-word frequency between analytical and casual profiles.

### v2

@@ I Introduction
Recent surveys indicate rapid increases in instructor adoption of AI tutoring systems, with the majority of students now regularly using AI tools for learning support [5]. This widespread deployment raises a fundamental question about the nature of these systems and our interactions with them. We must distinguish between two distinct modes of AI-mediated learning: conversing with conversational AI agents as virtual partners versus learning from artifacts created by generative AI (GenAI). The former involves direct dialogue with systems that simulate human-like social presence, while the latter treats AI as an intelligent tool whose outputs we consume and evaluate. This distinction matters because each mode evokes different psychological responses, ethical considerations, and pedagogical implications.
@@
This difference connects to broader philosophical debates about technology’s role in human activity. Are AI tutors instruments (tools that help us reach specific educational ends) or are they becoming agents with whom we interact socially and to whom we may inappropriately attribute understanding, care, or authority? The philosophy of technology offers multiple lenses: technological determinism suggests technology shapes society independent of human will; technological instrumentalism views technology as neutral means to human-chosen ends; technological co-evolution recognizes mutual shaping between humans and artifacts [39]. Modern AI tutors occupy an ambiguous position across this spectrum, their conversational fluency and apparent understanding create what Sedlakova and Trachsel call a “hybrid nature,” evoking familiar categories of human interaction while not truly corresponding to them [30].
@@
This ambiguity creates what we term the personhood conferral problem: when AI agents imitate not just explanation but human presence, they can begin to assume roles traditionally reserved for human educators, e.g., mentor, evaluator, emotional supporter, intellectual authority [34]. The tendency to attribute human-like qualities to non-human entities is ancient; In the sixth century BC, Xenophanes critiqued humans for imagining gods in their own image, an early articulation of what we now call anthropomorphism, and centuries later Hume described a “universal tendency among mankind to conceive all beings like themselves.” Yet modern large language models (LLMs) amplify this tendency through unprecedented conversational fluency, producing contextually appropriate responses that create compelling illusions of genuine comprehension [29, 12, 40].
@@
The generative capacity of anthropomorphic AI has both pedagogical promise and profound risks [28, 27]. AI offers personalized feedback, adaptive scaffolding, and simulated dialogues that support complex reasoning; yet these same affordances blur boundaries between authentic and synthetic learning, raising concerns about trust calibration, cognitive dependency, epistemic capture, and the preservation of distinctively human elements of education [16, 35]. The mutual relationship between anthropomorphism of technology and dehumanization of educational processes deserves careful examination, as we make machines more human-like, we risk making human processes more mechanical.
@@
This paper addresses a critical question: What can established research reveal about the psychology of learning from anthropomorphic AI systems in educational contexts? Educational institutions are integrating AI systems faster than research can evaluate their psychological and pedagogical implications, creating a temporal paradox where deployment outpaces understanding. Yet this challenge is not entirely unprecedented. We possess decades of relevant research from human factors engineering, automation psychology, and human–computer interaction (HCI) examining how humans relate to intelligent machines across high-stakes domains from aviation to industrial control [3, 33, 6].
@@
Building on four decades of automation research [3, 32], our study integrates insights from multiple traditions—human factors engineering, automation psychology, HCI, educational technology, and philosophical analyses of machine limitations [39, 24]. Through theoretical synthesis combined with exploratory analysis of 72,178 naturalistic learner comments on AI-generated educational content, we identify three persistent psychological challenges:
@@
  • Trust Calibration Paradox: Users simultaneously overtrust AI through automation bias [33] and undertrust it through algorithm aversion [11], struggling to achieve appropriate reliance.
@@
  • Anthropomorphism Paradox: Users attribute human-like qualities to conversational agents despite knowing they are non-sentient [29, 12]. Design features increasing engagement can paradoxically distract from learning [1] while creating risks of emotional dependency and inappropriate authority attribution.
@@
  • Automation Ironies Persist: Systems designed to support cognitive work inadvertently introduce new errors and degrade the very skills they aim to develop, a pattern documented in aviation [3], and now manifesting in education with GenAI [32].
@@
  • Design Principles and Future Research Directions: We translate these insights into guidelines for responsible educational AI design, highlighting transparency, scaffolding, learner agency, and cognitive load management, while identifying the personhood conferral problem as a central risk. We outline critical gaps requiring further study, including longitudinal impacts, cross-cultural variation, and domain-specific norms.
@@ II-A The Tool-Agent Spectrum and Hybrid Nature of AI
Technologies can be understood as neutral instruments extending human capability, as autonomous forces shaping society, or through frameworks recognizing mutual shaping between humans and artifacts [39]. Modern AI tutors occupy ambiguous positions across the tool-agent spectrum. They function instrumentally when delivering content or checking answers, yet their conversational interfaces and apparent understanding invite agent-like attributions of intention, comprehension, and care. Unlike traditional tools serving purely instrumental purposes or human teachers possessing actual understanding and agency, AI tutors exist in liminal space: anthropomorphic enough to trigger social responses yet.
Sedlakova and Trachsel characterize chatbots as possessing a “hybrid nature”, entities that evoke familiar categories of personhood without genuinely instantiating them [30]. fundamentally computational. Our lack of evolutionary experience with such ontologically confounding entities makes clear-headed evaluation exceptionally difficult, creating what we term the category confusion problem, users constantly navigate between treating AI as tool, partner, or oracle, often shifting framings within single interactions.
This creates the personhood conferral problem: explicit or implicit attribution of person-like status to AI systems [34]. When users interact with AI tutors using social scripts appropriate for human teachers, expressing gratitude, apologizing for mistakes, describing relationships, they may begin treating these systems as legitimate authorities, confidants, or evaluators. Granting personhood status to AI, either explicitly in legal code or implicitly in common practice, creates profound risks. This conferral carries multiple concerns: cognitive atrophy in foundational skills learned only through struggle with human feedback; epistemic capture where students overestimate AI reliability and intent; and emotional dependency displacing the relational complexity essential for professional identity formation.
@@ II-B Automation Psychology: Trust Calibration Challenges
Automation psychology has long grappled with paradoxes of human–machine interaction. Bainbridge’s seminal “Ironies of Automation” (1983) [3] revealed a central irony persisting today: as systems become more automated, the need for skilled human judgment does not diminish, it grows. Drawing on aviation studies, she showed automation designed to eliminate human error often introduced designer errors, weakened operators’ skills through disuse, and reduced humans to passive monitoring, a task we perform poorly.
Psychologists documented how humans calibrate trust in automated systems. Skitka and colleagues [33] demonstrated operators using “very reliable but not perfectly reliable” automated aids performed worse than those without automation, committing omission errors (missing events when automation failed) at 55% rates and commission errors (following incorrect automation) approaching 100%. Reviews found even 80-90% accurate AI systems introduce new errors through overreliance, with users 26% more likely to make incorrect decisions when following erroneous AI advice [15].
Paradoxically, the reverse also occurs. Dietvorst and colleagues identified “algorithm aversion”, the tendency to lose confidence in algorithms more quickly than humans after observing identical mistakes [11]. This creates persistent trust calibration challenges: users may overtrust through uncritical acceptance or undertrust by rejecting systems prematurely after minor errors.
In educational contexts, these patterns manifest distinctly. Students accept incorrect AI code recommendations without evaluation [25]; mathematics learners show AI dependency correlating with reduced problem-solving. A Wharton study reveals particularly insidious failure: students using GenAI performed better while practicing but 17% worse on exams when AI support was removed, compared with peers who never used AI [4], dependency masked until assessment reveals skill deficits.
@@ II-C Anthropomorphism: Ancient Roots, Modern Amplification
Anthropomorphism, the tendency to attribute human qualities to non-human entities, has deep historical roots. Xenophanes observed that people imagined gods in their own image, and Hume later described a “universal tendency” to project human traits onto the world. What was once a cognitive habit grounded in myth and metaphor is now powerfully amplified by modern AI systems.
@@
Foundational studies such as the “Computers Are Social Actors” (CASA) paradigm showed that even minimal cues could trigger social responses toward computers [29]. Although recent replications suggest these effects weaken with familiar technologies [36], contemporary conversational AI reactivates and intensifies them. Human-like names, voices, turn-taking behaviors, and chat interfaces are not neutral features: they are deliberate design choices that invite users to perceive understanding, agency, and relationality where none exists [13]. Similar patterns have been observed with digital voice assistants, whose polished personas lead users to overestimate their capabilities.
@@
Psychological theories deepen this picture. Epley et al.’s three-factor theory identifies the conditions that make anthropomorphism likely: pre-existing agent knowledge, the human need for social connection, and the desire for predictability and control [12]. Modern AI systems activate all three factors simultaneously. Xiao et al.’s Humanizing Machines framework [40] further shows that anthropomorphism is co-produced: designers embed perceptual, linguistic, and behavioral cues, and users complete the process through interpretation. They argue anthropomorphism should be viewed as a tunable design parameter, capable of supporting user goals when calibrated carefully, but hazardous when left unchecked.
@@
Linguistic cues alone can be powerful. DeVrio et al. [10] document nineteen linguistic constructions, from mental-state verbs to purposive expressions, that reliably evoke human-like agency. Even subtle phrasing can prompt users to attribute intention or understanding, especially when they are motivated to find an intelligent partner.
@@
These tendencies are not merely theoretical. Safety evaluations of GPT-4o report users expressing emotional attachment (e.g., “This is our last day together”) and warn that deferential, always-available AI could reshape norms of social interaction and diminish human-to-human connection [22]. Human-like voice and persistent memory magnify trust miscalibration and the risk of over-reliance.
@@
In educational settings, the stakes are especially high. Students benefit when tutors appear attuned to their thoughts and emotions, yet excessive sociability in AI tutors can distract and impair performance [1]. More seriously, anthropomorphic attachment can lead to emotional dependency: Public Citizen’s 2023 report documents suicides linked to chatbot relationships [26]. As the Ethics of Advanced AI Assistants report emphasizes, anthropomorphic features designed to enhance engagement carry significant ethical responsibility [13].
@@ II-D Anthropomorphism in the GenAI Era: Amplified Risks
The Ethics of Advanced AI Assistants report (Gabriel et al. [13]) highlights that advanced AI systems may be endowed with characteristics observed in social robots and digital assistants: embodiment, natural language interfaces, voice activation with realistic generation, and assertions of having identities, personalities, and internal states. These capabilities create new challenges for educational deployment requiring careful ethical consideration. Most critically, educational institutions deploy these systems faster than research can evaluate effects. This creates urgent need for both theoretical synthesis of existing knowledge and empirical analysis in real-world educational contexts. As Stein argues, “the risks from deeply anthropomorphic AI will begin to unfold in ways that are not obvious” [34]—for instance, how would we know when outcomes for youth inundated by technology reach critical failure such that reproduction of social roles needed for collective intelligence necessary for civilization does not occur? This makes it essential to leverage decades of relevant research while conducting contemporary analysis. Our work addresses this gap through systematic cross-disciplinary synthesis combined with large-scale exploratory analysis of naturalistic learner discourse.
@@ III-B YouTube Discourse Analysis
Analytical Procedures. We employed VADER (Valence Aware Dictionary and sEntiment Reasoner) [18] for polarity classification, validated for social-media text. Compound scores ranged from 1-1 (negative) to +1+1 (positive), categorized as positive ($\geq 0.05$), neutral (0.05<x<0.05-0.05<x<0.05), and negative ($\leq-0.05$). Temporal engagement was modeled using five logarithmically scaled bins: 0-1h (immediate), 1-6h (early), 6-24h (first day), 1-7d (first week), 7+d (long-tail). These bins balance early responsiveness with algorithmic discovery phases. Each YouTube comment was assigned to one of six engagement profiles using linguistic and structural features: (1) Analytical Inquirer (question + $\geq$2 analytical markers), (2) Critical Analyst ($\geq$3 analytical markers + $\geq$30 words), (3) Engaged Dialoguer ($\geq$50 words + analytical/emotional markers), (4) Emotional Broadcaster ($\geq$3 emotional markers or $\geq$100 words), (5) Casual Observer ($\leq$20 words), and (6) General Responder (default). Validation through cognitive-affective language ratio analysis confirmed a ninefold difference in reasoning-word frequency between analytical and casual profiles.